OpenStudy (anonymous):

Use f(x) = 5x − 2 and g(x) = 4 − x^2 to evaluate the expression (g o f)(x)

OpenStudy (anonymous):

(g o f)= 4 - (5x-2)^2 (g o f)= 4- (25x^2 -10x -10x + 4) = 4- (25x^2 - 20x + 4) simplify from there